OpenStudy (anonymous):

how can we say that 1+1/2+1/3+1/4...is not a convergent series

OpenStudy (anonymous):

is no one here to answer my questions???????????

OpenStudy (anonymous):

ok i m going bcz no one is ready to answer me

OpenStudy (anonymous):

not converging because this famous series, known as the harmonic series, is too dang big

OpenStudy (anonymous):

i bet you can find a good proof on line, but the basic idea is to go out far enough so that the last denominator is a power of 2, add and see that you get 1/2. for example \[\frac{1}{3}+\frac{1}{4}>\frac{1}{4}+\frac{1}{4}=\frac{1}{2}\] and so on

OpenStudy (anonymous):

\[\frac{1}{5}+\frac{1}{6}+\frac{1}{7}+\frac{1}{8}>4\times \frac{1}{8}=\frac{1}{2}\] etc etc